OpenStudy (anonymous):

find the value of log3(1/9) w/out using a calculator

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

3^x=1/9

OpenStudy (anonymous):

do you know how to do it with a calculator?

OpenStudy (slaaibak):

\[\log_3 (1/9) = \log_3(3^{-2}) = -2 * \log_3(3) = -2*1 = -2\]